OpenStudy (anonymous):

A bag contains 25 paise, 10 paise and 5 paise coins in the ratio 1 : 2 : 3. If the total value of the coins is Rs. 36, the number of 25 paise coins is

OpenStudy (anonymous):

see x+2x+3x=36 so,x=6 or 36/(1+2+3)=6,so number of 25 paise coins is 24 but the ans wer is 60?.

OpenStudy (anonymous):

25x/100+10*2x/100+5*3x/100=36 60x/100=36,so x=60

OpenStudy (anonymous):

how if rs6 is the value of 25 paise coins.the total coin value=36 bcoz,10 is rs12 and 18rs for 5 paise

OpenStudy (anonymous):

so the number of 25 coins is 24 only ?.25 paise to make one rupee is 4 so 4*6=24

ganeshie8 (ganeshie8):

lets start from scratch

ganeshie8 (ganeshie8):

say, there are "a" 25 paise coind, "b" 10 paise coins, and "c" 5 paise coins in the bag : 25a + 10b + 5c = 3600

ganeshie8 (ganeshie8):

fine with that equation ?

OpenStudy (anonymous):

yes so a=60 coins,120 for10,180 for5

OpenStudy (anonymous):

so there value bcomes 3600 right?

ganeshie8 (ganeshie8):

exactly !

OpenStudy (anonymous):

but they gave 36 only for total value.this doesnt match

ganeshie8 (ganeshie8):

36 rupees = 3600 paise

ganeshie8 (ganeshie8):

notice that we're measuring in paise on left hand side

ganeshie8 (ganeshie8):

so right hand side becomes 3600

ganeshie8 (ganeshie8):

if you want to measure in rupees, you would get : 0.25a + 0.1b + 0.05c = 36

ganeshie8 (ganeshie8):

because 25 paise = 0.25 of a rupee 10 paise = 0.1 of a rupee 5 paise = 0.05 of a rupee

ganeshie8 (ganeshie8):

in short : 1 paise = 0.01 of a rupee

ganeshie8 (ganeshie8):

or : 1 rupee = 100 paise so, 1/100 rupee = 1 paise which is same as : 0.01 rupee = 1 paise

ganeshie8 (ganeshie8):

all above are same ^^ just messing with the `paise <-> rupee` conversion

OpenStudy (anonymous):

if we get 24 coins of 25 paise this also will be correct?

OpenStudy (anonymous):

they asked for 25 paise coins so 24 only.not for one paise coin..it still confusing question.i understood it taken for paise

ganeshie8 (ganeshie8):

lets see : 25 paise : 24 coins 10 paise : 24*2 coins 5 paise : 24*3 coins

ganeshie8 (ganeshie8):

that means : 25 paise : 24 coins = 6 rupees 10 paise : 24*2 coins = 4.8 rupees 5 paise : 24*3 coins = 3.6 rupees -------------------------------------

ganeshie8 (ganeshie8):

you won't get a total of 36 rupees, so 25 paise coins can NOT be just 24 okay ?

ganeshie8 (ganeshie8):

25 paise coins have to be 60

OpenStudy (anonymous):

yes.but little twisted question

OpenStudy (anonymous):

got it!

ganeshie8 (ganeshie8):

agree, here is the complete solution in one place : say, there are "a" 25 paise coins, "b" 10 paise coins, and "c" 5 paise coins in the bag : 25a + 10b + 5c = 3600 ------(1) we're give n : a : b : c = 1 : 2 : 3 so b = 2a, c = 3a substitute these in equation (1), you get : 25a + 10(2a) + 5(3a) = 3600 60a = 3600 a = 60

OpenStudy (anonymous):

good:)